A garden has an area of 96 yd. Its width is 6 yd. Find its perimeter. How do I do that?
Found 3 solutions by chriswen, ewatrrr, unlockmath:
Answer by chriswen(106)

A=l*w
The area is 96yd and the width is 6yd.
96=l*6
96/6=l
l=16
The length of the garden is 16yd.
P=2(l+w)
P=2(16+6)
P=2(22)
P=44
The perimeter of the garden is 44yd.

Answer by ewatrrr(24785)

 
Hi

Let x and y represent the width and length respectively
P = 2*y + 28x
AND x*y = 96yd^2
6yd = 96yd^2/y |Its width is 6 yd.
y = 96yd^2/6yd = 16yd
2*y + 2*x = P
2*16yd + 2*6yd = 44 yd

Answer by unlockmath(1688)

Hello,
We know the area of a shape is width times height, right?
We can use that to form this equation (Let x be the height):
6x=96
Divide both sides by 6 to get:
x=16 yds is the height
I'll see if you can figure out the perimeter.
Make sense?
RJ
www.math-unlock.com